2. Asset in a the average c O 8.01 x O 12.98x equal market share with sales of $200,000 each. Youve collected company data to compare Like Games and Our Play. Last year, the average of these two (in d 7,700
media%2Fd42%2Fd424122f-8936-4099-9da9-e6
Answer #1

Answer 1:

Correct answer is:

11.80x

Explanation:

Monroe Manufacturing:

Annual sales = $200,000

Total Current assets = $56,500

Cash = $25,425

Account Receivable = $14,125

Hence inventory = Total current assets - Cash - Account receivable = 56500 - 25425 -14125 = $16,950

Inventory turnover = 200000 / 16950 = 11.80

Hence option D is correct and other options are incorrect.

Answer 2:

Correct answer (statement that is true) i:

Monroe Manufacturing is holding less inventory per dollar of sales compared to the industry average.

Explanation:

Monroe Manufacturing's Inventory turnover ratio = 11.80x

Industry average inventory turnover ratio = 10.03x

Hence the statement that Monroe Manufacturing is holding less inventory per dollar of sales compared to the industry average is true.

Answer 3:

Let us calculate the required ratios:

Like Games Our Pla Sales Account receivable (AR Net fixed assets (FA Total assets (TA $200,000$200,000 $7,800 $110,000 S160,000 $250,000 Industry average $510,000 $7,700 $433,500 $469,200 $5,400 $190,000 Days sales outstanding [AR/ (sales/365)] Fixed Asset Turnover [ Sales FA] Total Asset Turnover ratio [Sales/ TA] 10 1.82 1.05 14 1.25 0.80 6 1.18 1.09

1. Our Play has 14 days of sales tied up in receivables which is much higher than industry average. It takes Our Play higher time to collect cash from its customers that it takes Like Games.

2. Like Games fixed assets turnover ratio is higher than that of Our Play. This is because Like Games was formed eight years ago, so the acquisition cost of its fixed assets at historic values when the company bought its assets and has been depreciated since then. Assuming that fixed asset prices (net book values) rose over the past six years due to inflation, Our Play paid a higher amount for its fixed assets.

3. The average total assets turnover in the electronic industry is 1.09x which means that $1.09 of sales is being generated with every dollar of investment in assets. A higher total assets turnover ratio indicates greater efficiency. Both companies' total assets turnover ratios are lower than the industry average.